OpenStudy (anonymous):

I'm having a terrible time with solving linear equations by the addition/elimination method. Here is my problem: 3x+9y=12, 5x-4y=3. Can anyone help?

OpenStudy (amistre64):

i assume that 5x+4y?

OpenStudy (anonymous):

or 5x-4y :p

OpenStudy (anonymous):

no, these are the equations given in the question. I'm really lost now.

OpenStudy (amistre64):

multiply or divide to get the same "value" of a term, but opposites, in each equation

OpenStudy (amistre64):

these are the exact equations? no typos??? 3x+9y=12 and 5x=4y=3

OpenStudy (anonymous):

3x+9y=12, 5x-4y=3

OpenStudy (anonymous):

LOL

OpenStudy (amistre64):

divide the top by 3 and the bottom by -5

OpenStudy (amistre64):

x+3y =4 -x+4y/5= -3/5 then add together and solve for y

hartnn (hartnn):

instead of dividing,wouldn't multiplying by 4 and 9 be easy??then add..

OpenStudy (amistre64):

define "easy" ;)

hartnn (hartnn):

not involving fractions here...

OpenStudy (amistre64):

x+3y =4 -x+4y/5= -3/5 --------------- 19y/5 = 17/5 y = 17/19

OpenStudy (amistre64):

not sure why "not involving fractions" makes it easier ... unless you fear pie ;)

OpenStudy (anonymous):

OK-now I'm totally lost. This is what I'm doing: 3x+9y=12 5x-4y=3 I combine these two equations to equal 8x+5y=15 IS THIS RIGHT?

hartnn (hartnn):

lol! :P i thought 12-45 would be easier for her than 3+4/5....

OpenStudy (amistre64):

no, you have not eliminated any of the variables by that process

OpenStudy (anonymous):

I mean for a starting equation, I realize I have to solve for the variables

OpenStudy (amistre64):

you havent created a situation where one of the variables is removed; all you have done is create a line that has an infinite number of solutions

OpenStudy (anonymous):

OK, then can you explain the steps? Please go slow, and try not to laugh. I'm 55 taking Algebra for the 4th time and I'm determined to get this stuff.

OpenStudy (amistre64):

elimination, by its namesake alone, eliminates one of the variables in the process so that we are left with a single variable to play with. does that make sense?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

does it make sense that we can scale an equation and it still retains its inherent solutions; that only the form of it is altered?

OpenStudy (anonymous):

scale?

OpenStudy (amistre64):

to scale means: to multiply or divide by a constant

OpenStudy (anonymous):

you can't solve any equation unless it has only one variable

OpenStudy (amistre64):

vikrant, thats not quite accurate.

OpenStudy (anonymous):

Then we can "scale" an equation and still retain the solution. Only the form is altered. Is that right?

OpenStudy (anonymous):

if you multiply an equation with any number... it doesn't effects... just try to make coefficients of any one variable equal.

OpenStudy (anonymous):

P.S-My English is poor

OpenStudy (anonymous):

I dont understand what you are saying

OpenStudy (amistre64):

correct so we want to scale the given equations so that we can eliminate one of the variables and only focus on a single variable

OpenStudy (anonymous):

OK

OpenStudy (amistre64):

for example y = x is a line equation if we scale this by 2 we get: 2y = 2x which has the same solutions as y=x if we scale this by -4 we get: -4y = -4x which has the same solutions as y=x if we scale this by any constant "c" we get: cy = cx which has the same solutions as y=x

OpenStudy (anonymous):

^^YAy!!

OpenStudy (anonymous):

got it

OpenStudy (amistre64):

so, lets look at your system of equations, and see what we can scale them by to help us eliminate one of the variables in the process

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

This is always where I get stuck.

OpenStudy (amistre64):

im going to focus on eliminating the x parts 3x+9y=12 5x-4y=3 if we had x up top and -x on the bottom; then x-x=0 correct?

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

so, 3x/3 = x 5x/-5 = -x so my simplest idea is to scale the top by 1/3 and the bottom by -1/5

OpenStudy (anonymous):

I don't understand where you got -5 from

OpenStudy (anonymous):

nevermind

OpenStudy (amistre64):

\[\frac{3x+9y=12}{3}\]\[\frac{5x-4y=3}{-5}\] \[x+3y=4\]\[-x+\frac{4}{5}y=-\frac35\]

OpenStudy (anonymous):

WHEW! I'm going to try and digest this now. I think with your help on this problem, I will be able to do more problems like it.

OpenStudy (amistre64):

keep in mind that we can scale these by ANY constant factor; so we did not have to use 1/3 and -1/5; those were just what I chose for convience to me

OpenStudy (anonymous):

so I can use ANY numbers?

OpenStudy (amistre64):

yes; but in the end you desire a situation where you can eliminate one of the variable

OpenStudy (anonymous):

So how did you come up with 1/3 and -1/5?

OpenStudy (anonymous):

because they make coefficients of x in both equations same..

OpenStudy (amistre64):

i wanted the situation where: x + (- x) = 0 (3x) + (5x) = 0 this tells me that i would want to divide the top by 3 and the bottom by -5

OpenStudy (anonymous):

ok-I think my light bulb is beginning to glow. You have been a great deal of help to me and I appreciate it beyond words. God Bless You for helping a poor old grandma get through this!

OpenStudy (amistre64):

we can use the same concept; and instead of the situation: x + (-x) = 0, we might think it easier to have 15x + (-15x) = 0 (3x) + (5x) = 0 then we could multiply the top by 5 and the bottom by -3, right? 5(3x) + (-3)(5x) = 0

OpenStudy (anonymous):

got to think on this...

OpenStudy (amistre64):

good luck with it ;)

OpenStudy (anonymous):

You are so very kind, I see the 15x +(-15x) = 0 now. I will continue to work on this, thanks again, you are my hero today! Deb

OpenStudy (phi):

Here is an example being worked out http://www.khanacademy.org/math/algebra/systems-of-eq-and-ineq/v/addition-elimination-method-2